# An Efficient Local Search Approach for Polarized Community Discovery in Signed Networks

### v1

@@ Appendix A Proofs
$c_{\text{sim}}\triangleq\sum_{(i,j)\in E}A_{i,j}.$ (14)
$c_{\text{abs}}\triangleq\sum_{(i,j)\in E}|A_{i,j}|.$ (15)
@@
$f^{\text{full}}$ $\triangleq N^{+}_{\text{intra}}-N^{-}_{\text{intra}}+N^{-}_{\text{inter}}-N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}-\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}$ (16)
$f^{\text{MaxAgree}}$ $\triangleq N^{+}_{\text{intra}}+N^{-}_{\text{inter}}$
@@
$f^{\text{MinDisagree}}$ $\triangleq N^{-}_{\text{intra}}+N^{+}_{\text{inter}}$
@@
$f^{\text{MaxCorr}}$ $\triangleq N^{+}_{\text{intra}}-N^{-}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}$
$f^{\text{MinCut}}$ $\triangleq-N^{-}_{\text{inter}}+N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}$
@@
$S^{(1)}=\{\{1,2,3\}\},\quad S^{(2)}=\{\{1,2\},\,\{3\}\},\quad S^{(3)}=\{\{1,3\},\,\{2\}\},\quad S^{(4)}=\{\{2,3\},\,\{1\}\},\quad S^{(5)}=\{\{1\},\,\{2\},\,\{3\}\}.$
@@ 14 $-\,N_{\text{intra}}^{-}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$1 0 -3 0 -1
For $k=2$, the full objective always increases (or remains constant) when an object is made non-neutral, aligning it with max agreement. To see this, consider a clustering with $k=2$ non-neutral clusters, and let $M_{i,m}=\sum_{j\in S_{m}}A_{i,j}$ be the total similarity of object $i$ to cluster $m$. The impact on the objective when assigning $i$ to $m$ is $M_{i,m}-M_{i,p}$, where $p$ is the other cluster. Since this difference is always positive when $i$ is placed in its most similar cluster, assigning $i$ as non-neutral always improves the objective. Then, since all objects are non-neutral, the problem is equivalent to the $k$-CC problem where we know max agreement and the full objective are equivalent (from Proposition 1). For $k>2$, this reasoning no longer holds, as contributions from other clusters can outweigh the within-cluster similarity to the most similar cluster (i.e., making the total contribution negative), potentially making neutrality optimal. However, we note that in our final objective (Eq. 7), when $\alpha$ and $\beta$ are involved, all terms will contribute with unique information even for $k=2$.
See 1
Proof.
$(N^{+}_{\text{intra}}-N^{-}_{\text{intra}})+\alpha(N^{-}_{\text{inter}}-N^{+}_{\text{inter}})-\beta\sum_{m\in[k]}|S_{m}|^{2}$ $=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}-\beta\sum_{m\in[k]}|S_{m}|^{2}$ (18)
The second equality follows from Proposition 3. Defining $c_{\text{sim}}\triangleq\sum_{(i,j)\in E^{\prime}}A_{i,j}$, we obtain:
Substituting this into Eq. 18 and simplifying:
$\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}$ $=\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha(c_{\text{sim}}-\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j})$ (20)
Defining $A^{\prime}_{i,j}=(1+\alpha)(A_{i,j}-\beta)$, we observe that since $c_{\text{sim}}$ is a constant across clustering solutions, the problem reduces to finding a partition of $V^{\prime}$ that maximizes $\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{\prime}_{i,j}$. This is equivalent to the max correlation objective (Eq. 4) applied to the transformed adjacency matrix $A^{\prime}$. By Proposition 1, this objective is equivalent to the $k$-CC problem (Problem 1). Thus, solving the $k$-PCD problem requires solving the $k$-CC problem on the instance $G^{\prime}=(V^{\prime},E^{\prime})$, meaning $k$-PCD is at least as hard as $k$-CC. Since correlation clustering is NP-hard [Bansal et al., 2004, Giotis and Guruswami, 2006], we conclude that $k$-PCD is also NP-hard. ∎
See 3
Proof.
We have
The second equality (line 3) holds because the number of pairs of objects inside cluster $m$ is $|S_{m}|^{2}$. A similar regularization is established in [Chehreghani, 2023] for the minimum cut objective, where it is shown that optimizing this minimum cut objective regularized with $-\beta\sum_{m\in[k]}|S_{m}|^{2}$ is equivalent to optimizing the max correlation objective (Eq. 4) with similarities shifted by $\beta$. However, their result specifically considers the full network partitioning of unsigned networks, where the initial pairwise similarities are assumed non-negative. Moreover, they use a different regularization in practice: they shift the pairwise similarities so that the sum of the rows and columns of the similarity matrix becomes zero. ∎
See 2
Proof.
$c_{im}\triangleq-\beta+2\sum_{j\in[n]\setminus\{i\}}\big{(}x_{jm}(A_{i,j}-\beta)-\alpha\sum_{p\in[k]\setminus\{m\}}x_{jp}A_{i,j}\big{)},\quad\text{for }m\in[k],\quad c_{i0}\triangleq 0.$ (24)
Then, we obtain
$\nabla_{i}f({\bm{x}}_{[n]})={\bm{c}}_{i}.$ (26)
$=(({\bm{x}}_{i}^{\ast})^{T}{\bm{c}}_{i}^{(t)}+C)-(({\bm{x}}_{i}^{(t)})^{T}{\bm{c}}_{i}^{(t)}+C)$
$=(({\bm{x}}_{i}^{\ast})^{T}-({\bm{x}}_{i}^{(t)})^{T}){\bm{c}}_{i}^{(t)}$
$=({\bm{x}}_{i}^{\ast}-{\bm{x}}_{i}^{(t)})\cdot\nabla_{i}f({\bm{x}}^{(t)}_{[n]})$
See 3
Proof.
$g({\bm{x}}_{[n]})\triangleq\max_{{\bm{s}}_{i}\in\Delta^{k+1},\forall i\in[n]}({\bm{s}}_{[n]}-{\bm{x}}_{[n]})\cdot\nabla f({\bm{x}}_{[n]}).$ (29)
Then, we recall that
$\tilde{g}_{t}=\min_{0\leq l\leq t-1}g({\bm{x}}^{(l)}_{[n]})$ (30)
is the smallest duality gap observed in Alg. 1 up until step $t$. As established by [Lacoste-Julien et al., 2013] for general domains, the FW duality gap can be decomposed as follows.
$g({\bm{x}}_{[n]})$ $\triangleq\max_{{\bm{s}}_{i}\in\Delta^{k+1},\forall i\in[n]}({\bm{s}}_{[n]}-{\bm{x}}_{[n]})\cdot\nabla f({\bm{x}}_{[n]})$ (31)
$=\sum_{i\in[n]}\underbrace{\max_{{\bm{s}}_{i}\in\Delta^{k+1}}({\bm{s}}_{i}-{\bm{x}}_{i})\cdot\nabla_{i}f({\bm{x}}_{[n]})}_{\triangleq g_{i}({\bm{x}}_{[n]})}$
Let $g_{i}({\bm{x}}_{[n]})\triangleq\max_{{\bm{s}}_{i}\in\Delta^{k+1}}({\bm{s}}_{i}-{\bm{x}}_{i})\cdot\nabla_{i}f({\bm{x}}_{[n]})$ be the duality gap related to block $i$. We have that the FW duality gap is the sum of the gaps from each block: $g({\bm{x}}_{[n]})=\sum_{i\in[n]}g_{i}({\bm{x}}_{[n]})$.
From Definition 2 (convergence rate), in order to prove the stated convergence rate, we need to show that $\mathbb{E}[\tilde{g}_{t}]\leq nh_{0}/t$. The structure of our proof is similar to the proof of Theorem 2 in [Thiel et al., 2019]. However, here we adapt it to our problem and make the proof more rigorous (including correction of a mistake in the proof by [Thiel et al., 2019]). A key difference is that our objective in Eq. 10 is not multilinear in the blocks $i$. Then, as shown in Eq. 22 of Theorem 2, the first quadratic term can be transformed into a linear one by assuming a discrete solution (which we showed holds at every step $t$).
$=\sum_{i\in[n]}\frac{1}{n}g_{i}({\bm{x}}^{(t)}_{[n]})$
$=\frac{1}{n}g({\bm{x}}_{[n]}^{(t)}).$
[…]
$=\mathbb{E}[g_{i}({\bm{x}}_{[n]}^{(t)})],$
where the last equality follows from the Law of Total Expectation (i.e., that $\mathbb{E}_{Y}[\mathbb{E}_{X}{[X|Y]}]=\mathbb{E}_{X}{[X}]$, where $X$ and $Y$ are random variables). We therefore have that $\frac{1}{n}\mathbb{E}[g({\bm{x}}_{[n]}^{(t)})]=\mathbb{E}[g_{i}({\bm{x}}_{[n]}^{(t)})]$, where the expectation is w.r.t. all randomly chosen blocks $i$ before step $t$. Now, from Theorem 2 we have that our objective satisfies
Then, we have
$=\sum_{t=0}^{T-1}\mathbb{E}[f({\bm{x}}^{(t+1)}_{[n]})-f({\bm{x}}^{(t)}_{[n]})]$
$=\mathbb{E}[f({\bm{x}}^{(T)}_{[n]})]-f({\bm{x}}^{(0)}_{[n]})$
$\leq OPT-f({\bm{x}}_{[n]}^{(0)}),$
$\frac{T}{n}\mathbb{E}[\tilde{g}_{T}]$ $\leq OPT-f({\bm{x}}^{(0)})$ (37)
$\Rightarrow\mathbb{E}[\tilde{g}_{T}]$ $\leq\frac{n(OPT-f({\bm{x}}^{(0)}))}{T}.$
$\mathbb{E}[\tilde{g}_{T}]\leq\frac{n\sum_{(i,j)\in E}|A_{i,j}|}{T}.$ (38)
which we aimed to show since it holds for any $T$. ∎
See 1
Proof.
From part (a) of Theorem 2, the current solution, ${\bm{x}}^{(t)}_{[n]}$, remains discrete (i.e., hard cluster assignments) at every step of Alg. 1 for all $i\in[n]$. Moreover, each step of Alg. 1 consists of placing object $i$ in the cluster $m\in\{0,\dots,k\}$ with maximal gradient $G_{i,m}$. By part (b) of Theorem 2, this is equivalent to placing object $i$ in the cluster that maximally improves our objective in Eq. 7. ∎
See 4
Proof.
From Theorem 2, we recall that since the current solution ${\bm{x}}_{[n]}^{(t)}$ always remains discrete, our objective can be written as
Therefore, we have
$G_{i,m}\triangleq[\nabla_{i}f({\bm{x}}^{(t)}_{[n]})]_{m}=c_{im},\quad\text{for}m\in[k].$ (41)
$G_{i,0}\triangleq[\nabla_{i}f({\bm{x}}^{(t)}_{[n]})]_{0}=0.$ (42)
Now, recall the defintions $M_{i,m}\triangleq 2\sum_{j\in S_{m}}A_{i,j}$, $\eta_{i}\triangleq\sum_{m\in[k]}M_{i,m}$ and $\beta_{im}=2\beta|S_{m}|-2\beta\bm{1}_{[i\in S_{m}]}$. Then, we note that $2\sum_{j\notin S_{0}}A_{i,j}=\sum_{m\in[k]}M_{i,m}$ (sum of all similarities from object $i$ to all non-neutral objects). Then, we obtain the final expression, i.e.,
$c_{im}=-\beta+(1+\alpha)M_{i,m}-\beta_{im}-\alpha\eta_{i},\quad\text{for }m\in[k],$ (45)
See 4

### v2

@@ Appendix B Proofs
The second equality follows from Prop. 1. Defining $c_{\text{sim}}\coloneqq\sum_{(i,j)\in E^{\prime}}A_{i,j}$, we obtain:
Substituting this into Eq. 13 and simplifying:
$\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}=$ (15)
$=\sum_{m\in[k]}\sum_{i,j\in S_{m}}(A_{i,j}-\beta)-\alpha(c_{\text{sim}}-\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j})$
Defining $A^{\prime}_{i,j}=(1+\alpha)(A_{i,j}-\beta)$, we observe that since $c_{\text{sim}}$ is a constant across clustering solutions, the problem reduces to finding a partition of $V^{\prime}$ that maximizes $\sum_{m\in[k]}\sum_{i,j\in S_{m}}A^{\prime}_{i,j}$. This is equivalent to the max correlation objective (Eq. 11) applied to the transformed adjacency matrix $A^{\prime}$. By Prop. 2, this objective is equivalent to the $k$-CC problem (Problem 1). Thus, solving the $k$-PCD problem requires solving the $k$-CC problem on the instance $G^{\prime}=(V^{\prime},E^{\prime})$, meaning $k$-PCD is at least as hard as $k$-CC. Since correlation clustering is NP-hard [5, 20], we conclude that $k$-PCD is also NP-hard. ∎
See 1
Proof.
We have
The second equality (line 3) holds because the number of pairs of objects inside cluster $m$ is $|S_{m}|^{2}$. A similar regularization is established in [10] for the minimum cut objective, where it is shown that optimizing this minimum cut objective regularized with $-\beta\sum_{m\in[k]}|S_{m}|^{2}$ is equivalent to optimizing the max correlation objective (Eq. 11) with similarities shifted by $\beta$. However, their result specifically considers the full network partitioning of unsigned networks, where the initial pairwise similarities are assumed non-negative. Moreover, they use a different regularization in practice: they shift the pairwise similarities so that the sum of the rows and columns of the similarity matrix becomes zero. ∎
See 2
Proof.
$c_{im}\coloneqq-\beta+2\sum_{j\in[n]\setminus\{i\}}\big{(}x_{jm}(A_{i,j}-\beta)-\alpha\sum_{p\in[k]\setminus\{m\}}x_{jp}A_{i,j}\big{)},\quad\text{for }m\in[k],\quad c_{i0}\coloneqq 0.$ (19)
Then, we obtain
$\nabla_{i}f({\bm{x}}_{[n]})={\bm{c}}_{i}.$ (21)
$=(({\bm{x}}_{i}^{\ast})^{T}{\bm{c}}_{i}^{(t)}+C)-(({\bm{x}}_{i}^{(t)})^{T}{\bm{c}}_{i}^{(t)}+C)$
$=(({\bm{x}}_{i}^{\ast})^{T}-({\bm{x}}_{i}^{(t)})^{T}){\bm{c}}_{i}^{(t)}$
$=({\bm{x}}_{i}^{\ast}-{\bm{x}}_{i}^{(t)})\cdot\nabla_{i}f({\bm{x}}^{(t)}_{[n]})$
See 3
Proof.
$g({\bm{x}}_{[n]})\coloneqq\max_{{\bm{s}}_{i}\in\Delta^{k+1},\forall i\in[n]}({\bm{s}}_{[n]}-{\bm{x}}_{[n]})\cdot\nabla f({\bm{x}}_{[n]}).$ (24)
Then, we recall that
$\tilde{g}_{t}=\min_{0\leq l\leq t-1}g({\bm{x}}^{(l)}_{[n]})$ (25)
is the smallest duality gap observed in Alg. 1 up until step $t$. As established by [29] for general domains, the FW duality gap can be decomposed as follows.
$g({\bm{x}}_{[n]})$ $\coloneqq\max_{{\bm{s}}_{i}\in\Delta^{k+1},\forall i\in[n]}({\bm{s}}_{[n]}-{\bm{x}}_{[n]})\cdot\nabla f({\bm{x}}_{[n]})$ (26)
$=\sum_{i\in[n]}\underbrace{\max_{{\bm{s}}_{i}\in\Delta^{k+1}}({\bm{s}}_{i}-{\bm{x}}_{i})\cdot\nabla_{i}f({\bm{x}}_{[n]})}_{\coloneqq g_{i}({\bm{x}}_{[n]})}$
Let $g_{i}({\bm{x}}_{[n]})\coloneqq\max_{{\bm{s}}_{i}\in\Delta^{k+1}}({\bm{s}}_{i}-{\bm{x}}_{i})\cdot\nabla_{i}f({\bm{x}}_{[n]})$ be the duality gap related to block $i$. We have that the FW duality gap is the sum of the gaps from each block: $g({\bm{x}}_{[n]})=\sum_{i\in[n]}g_{i}({\bm{x}}_{[n]})$.
From Definition 2 (convergence rate), in order to prove the stated convergence rate, we need to show that $\mathbb{E}[\tilde{g}_{t}]\leq nh_{0}/t$. The structure of our proof is similar to the proof of Theorem 2 in [43]. However, here we adapt it to our problem and make the proof more rigorous (including correction of a mistake in the proof by [43]). A key difference is that our objective in Eq. 5 is not multilinear in the blocks $i$. Then, as shown in Eq. 17 of Thm. 2, the first quadratic term can be transformed into a linear one by assuming a discrete solution (which we showed holds at every step $t$).
$=\sum_{i\in[n]}\frac{1}{n}g_{i}({\bm{x}}^{(t)}_{[n]})$
$=\frac{1}{n}g({\bm{x}}_{[n]}^{(t)}).$
[…]
$=\mathbb{E}[g_{i}({\bm{x}}_{[n]}^{(t)})],$
where the last equality follows from the Law of Total Expectation (i.e., that $\mathbb{E}_{Y}[\mathbb{E}_{X}{[X|Y]}]=\mathbb{E}_{X}{[X}]$, where $X$ and $Y$ are random variables). We therefore have that $\frac{1}{n}\mathbb{E}[g({\bm{x}}_{[n]}^{(t)})]=\mathbb{E}[g_{i}({\bm{x}}_{[n]}^{(t)})]$, where the expectation is w.r.t. all randomly chosen blocks $i$ before step $t$. Now, from Thm. 2 we have that our objective satisfies
Then, we have
$=\sum_{t=0}^{T-1}\mathbb{E}[f({\bm{x}}^{(t+1)}_{[n]})-f({\bm{x}}^{(t)}_{[n]})]$
$=\mathbb{E}[f({\bm{x}}^{(T)}_{[n]})]-f({\bm{x}}^{(0)}_{[n]})$
$\leq OPT-f({\bm{x}}_{[n]}^{(0)}),$
$\frac{T}{n}\mathbb{E}[\tilde{g}_{T}]$ $\leq OPT-f({\bm{x}}^{(0)})$ (32)
$\Rightarrow\mathbb{E}[\tilde{g}_{T}]$ $\leq\frac{n(OPT-f({\bm{x}}^{(0)}))}{T}.$
$\mathbb{E}[\tilde{g}_{T}]\leq\frac{n\sum_{(i,j)\in E}|A_{i,j}|}{T}.$ (33)
which we aimed to show since it holds for any $T$. ∎
See 1
Proof.
From part (a) of Thm. 2, the current solution, ${\bm{x}}^{(t)}_{[n]}$, remains discrete (i.e., hard cluster assignments) at every step of Alg. 1 for all $i\in[n]$. Moreover, each step of Alg. 1 consists of placing object $i$ in the cluster $m\in\{0,\dots,k\}$ with maximal gradient $G_{i,m}$. By part (b) of Thm. 2, this is equivalent to placing object $i$ in the cluster that maximally improves our objective in Eq. 3. ∎
See 4
Proof.
From Thm. 2, we recall that since the current solution ${\bm{x}}_{[n]}^{(t)}$ always remains discrete, our objective can be written as
Therefore, we have
$G_{i,m}\coloneqq[\nabla_{i}f({\bm{x}}^{(t)}_{[n]})]_{m}=c_{im},\quad\text{for }m\in[k].$ (36)
$G_{i,0}\coloneqq[\nabla_{i}f({\bm{x}}^{(t)}_{[n]})]_{0}=0.$ (37)
Now, recall the defintions $M_{i,m}\coloneqq 2\sum_{j\in S_{m}}A_{i,j}$, $\phi_{i}\coloneqq\sum_{m\in[k]}M_{i,m}$ and $\beta_{im}=2\beta|S_{m}|-2\beta\bm{1}_{[i\in S_{m}]}$. Then, we note that $2\sum_{j\notin S_{0}}A_{i,j}=\sum_{m\in[k]}M_{i,m}$ (sum of all similarities from object $i$ to all non-neutral objects). Then, we obtain the final expression, i.e.,
$c_{im}=-\beta+(1+\alpha)M_{i,m}-\beta_{im}-\alpha\phi_{i},\quad\text{for }m\in[k],$ (40)
See 2
Proof.
@@
$c_{\text{sim}}\coloneqq\sum_{(i,j)\in E}A_{i,j}.$ (41)
$c_{\text{abs}}\coloneqq\sum_{(i,j)\in E}|A_{i,j}|.$ (42)
@@
$f^{\text{full}}$ $\coloneqq N^{+}_{\text{intra}}-N^{-}_{\text{intra}}+N^{-}_{\text{inter}}-N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}-\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}$ (43)
$f^{\text{MaxAgree}}$ $\coloneqq N^{+}_{\text{intra}}+N^{-}_{\text{inter}}$
@@
$f^{\text{MinDisagree}}$ $\coloneqq N^{-}_{\text{intra}}+N^{+}_{\text{inter}}$
@@
$f^{\text{MaxCorr}}$ $\coloneqq N^{+}_{\text{intra}}-N^{-}_{\text{intra}}=\sum_{m\in[k]}\sum_{i,j\in S_{m}}A_{i,j}$
$f^{\text{MinCut}}$ $\coloneqq-N^{-}_{\text{inter}}+N^{+}_{\text{inter}}=\sum_{m\in[k]}\sum_{p\in[k]\setminus\{m\}}\sum_{\begin{subarray}{c}i\in S_{m}\\ j\in S_{p}\end{subarray}}A_{i,j}$
@@
$S^{(1)}=\{\{1,2,3\}\},\quad S^{(2)}=\{\{1,2\},\,\{3\}\},$
$S^{(3)}=\{\{1,3\},\,\{2\}\},\quad S^{(4)}=\{\{2,3\},\,\{1\}\},\quad S^{(5)}=\{\{1\},\,\{2\},\,\{3\}\}.$
@@ 14 $-\,N_{\text{intra}}^{-}+N_{\text{inter}}^{-}-N_{\text{inter}}^{+}$1 0 -3 0 -1
For $k=2$, the full objective always increases (or remains constant) when an object is made non-neutral, aligning it with max agreement. To see this, consider a clustering with $k=2$ non-neutral clusters, and let $M_{i,m}=\sum_{j\in S_{m}}A_{i,j}$ be the total similarity of object $i$ to cluster $m$. The impact on the objective when assigning $i$ to $m$ is $M_{i,m}-M_{i,p}$, where $p$ is the other cluster. Since this difference is always positive when $i$ is placed in its most similar cluster, assigning $i$ as non-neutral always improves the objective. Then, since all objects are non-neutral, the problem is equivalent to the $k$-CC problem where we know max agreement and the full objective are equivalent (from Prop. 2). For $k>2$, this reasoning no longer holds, as contributions from other clusters can outweigh the within-cluster similarity to the most similar cluster (i.e., making the total contribution negative), potentially making neutrality optimal. However, we note that in our final objective (Eq. 3), when $\alpha$ and $\beta$ are involved, all terms will contribute with unique information even for $k=2$.
Appendix C Impact of $\alpha$ and $\beta$
In this section, we analyze the impact of $\alpha$ and $\beta$ in Eq. 3. In Appendix G.7 we investigate their impact experimentally. We begin by stating the following proposition.
Proposition 4.